# Centro de Física do Porto e Departamento de Física e Astronomia, Faculdade de Ciências da Universidade do Porto, Rua do Campo Alegre 687, 4169-007, Porto, Portugal τCPHT, CNRS, École Polytechnique, Institut Polytechnique de Paris, Palaiseau, France

### v1

@@ 2.1 Numerical analysis
$H=\sum_{i}H^{(1)}_{i}+\sum_{i}H^{(2)}_{i}-K\sum_{i}H^{(1)}_{i}H^{(2)}_{i}\,,$ (20)
which manifestly preserves two copies of the Fibonacci symmetries, as well as the diagonal spin flip $\mathbb{Z}_{2}$, since the deforming operators is at zero momentum and a permutation $\mathbb{Z}_{2}$ swapping the two chains. Indeed, the deformation
[…]
@@
The coefficients in front of the continuum operators depend on the details of the lattice model and are unfixed. We can study the two-coupled chain model on a closed chain of size $L$ and measure the entanglement entropy of a segment with size $l$. The result at $K/J=\infty$ is given in Figure 5. Picking the data points with $\log(l_{\rm conformal})>3.25$ and fitting against equation (7), we get the central charge
$c=1.151\pm 0.001.$ (23)
We emphasize here that the error reported above is the standard error of the linear-fit parameter. The central charge of the CFT can be (slightly) smaller than the above value due to finite size corrections.
@@
We can also measure the effective central charge as a function of $K/J$ to study the phase diagram. We first measure the entanglement entropy as a function of $l$ and then fit the against (7) to get $c_{\rm eff}$, see Figure 6. We find that at zero coupling, we get close to the decoupled value for the central charge $c\approx 1.4$, while for negative coupling, $c_{\rm{eff}}$ quickly drops to zero, signaling a gapped phase. Finally, for positive coupling, the effective central charge approaches a constant value, indicating a critical phase.
Figure 6: The effective central charge as a function of $K/J$. This is a measurement at $L=80$.
In addition to the entanglement entropy, we can measure the ground state energy as a function of the lattice size, as given in Figure 7.
@@
[…]
2.1.1 Phase diagram and energy levels
Combining our results with various earlier works, we now have a fairly complete understanding of $\mathcal{C}$-preserving RG flows from the decoupled tricritical Ising fixed point ($2\times\mathcal{M}_{5,4}$). The flows are summarized in Figure 8. Firstly, the ($2\times\mathcal{M}_{5,4}$) fixed point has one relevant direction towards the gapped phase. In fact, the pure $\sigma^{\prime(1)}\sigma^{\prime(2)}$ deformation leads to a gapped phase which was described in LECLAIR1998523 . This flow is integrable, and the S-matrix of its massive excitations can be determined exactly. Then, the $\epsilon^{\prime\prime}\equiv\phi_{(3,1)}$ deformation is the leading irrelevant operator that controls the (inverted) Zamolodchikov flows, i.e. we expect to be able to reach a decoupled tetracritical Ising fixed point ($2\times\mathcal{M}_{6,5}$) which is fully unstable. [^6] The $\sigma^{\prime(1)}\sigma^{\prime(2)}$ deformation of ($2\times\mathcal{M}_{6,5}$) also leads to a massive phase, this is again discussed in LECLAIR1998523 . The fixed point we have found with $c\approx 1.15$ is of course stable under $\mathcal{C}$ preserving deformations and we will argue that it is irrational below. For consistency with the topology of the phase diagram, a fourth fixed point, with one stable and one unstable direction, is needed. This is denoted as CFT in Figure 8. We will propose a candidate for this fixed point in the next section as well. [^7] It should be possible to reach this unstable fixed point by tuning a second $\mathcal{C}$ invariant coupling in the lattice model. We leave this analysis for future work.
Figure 8: RG flow of the 2-chain model. The blue arrow corresponds to the $K/J$. Notice the plot is symmetrical along the vertical axis.
Having understood how our fixed point fits in the phase diagram, we can also list out a few excited states and therefore determine low-lying operator dimensions. We have already measured the central charge $c$ in Figure 5. The result in Figure 7 then allows us to measure $E_{\infty}$ and the effective “speed of light” $v$ that appears in equation (6). Together they give us the relation between the energy spectrum and the scaling dimension of operators. They are summarized in Figure 9.
@@ 2.1.1 Phase diagram and energy levels
Figure 9: The spectrum of the coupled 2-chain CFT measured at $K/J=\infty$. The size of the system is $L=40$. The system has a $\mathbb{Z}_{2}^{(p)}$ permutation symmetry which permutates the states on the two chains. The red dots/blue circles indicate states that are even/odd under the $\mathbb{Z}_{2}^{(p)}$ symmetry. In addition to $\mathbb{Z}_{2}^{(p)}$, the system has an emergent $\mathbb{Z}_{2}$ symmetry, which corresponds to translating the lattice by one site. This can be seen from the fact that the low-energy states cluster around $p=0$ and $p=\pi$.
Remarkably, we find several scalar states that are doubly degenerate. This cannot be explained from the $\mathbb{Z}_{2}^{(p)}\times\mathbb{Z}_{2}$ symmetry alone, since this is an abelian group and therefore has only one-dimensional representations. These degeneracies are obvious in the decoupled UV theory. Given any two operators $\mathcal{O}$ and $\mathcal{O}^{\prime}$ in a given copy of tricritical Ising, we can always use them to build two degenerate states in the tensor product theory:
$\mathcal{C}\subset(\rm{Fibonacci}\times Fibonacci)\rtimes\mathbb{Z}_{2}\,,$ (25)
in full analogy with the existence of two dimensional representations for the invertible non-abelian symmetry $D_{4}=(\mathbb{Z}_{2}\times\mathbb{Z}_{2})\rtimes\mathbb{Z}_{2}$, where the rightmost $\mathbb{Z}_{2}$ acts by permuting the actions of the first two. [^8] Consider the two-dimensional Hilbert space spanned $|\sigma\rangle\otimes|1\rangle$ and $|1\rangle\otimes|\sigma\rangle$, the symmetry acts on it as $W^{(1)}=\left(\begin{array}[]{cc}-\varphi^{-1}&0\\ 0&\varphi\\ \end{array}\right),\quad W^{(2)}=\left(\begin{array}[]{cc}\varphi&0\\ 0&-\varphi^{-1}\\ \end{array}\right),\quad\mathbb{Z}_{2}^{(p)}=\left(\begin{array}[]{cc}0&1\\ 1&0\\ \end{array}\right).$ (32) One can easily see that the two-dimensional Hilbert space does not have an invariant subspace under the action of the above matrices. The two states, therefore, should have the same energy. The precise mathematical statement is related to the G-extension of modular tensor categories, see for example delaney2019fusionrulespermutationextensions . We will leave a rigorous study of this symmetry to future work.
2.2 Continuum limit and candidate fixed points
In the continuum limit, we can describe our system by the following formal action
$S=\sum_{i=1}^{2}S_{(i)}^{\textrm{tricrit}}+g\int d^{2}x\,\sigma^{\prime}_{(1)}\sigma^{\prime}_{(2)}+\rm{irrelevant}\,.$ (33)
The invertible symmetry of the model is $\mathbb{Z}^{(p)}_{2}\times\mathbb{Z}_{2}$, coming from the diagonal sign-flip $\sigma_{(i)}\to-\sigma_{(i)}$ and the permutation $\mathbb{Z}^{(p)}_{2}$ $\sigma_{(1)}\leftrightarrow\sigma_{(2)}$. Additionally we preserve two copies of the non-invertible Fibonacci lines.
Based on the relation to the $\mathbb{Z}_{8}$ parafermion theory and the RG flows studied in Fateev:1991bv , the authors of Chang:2018iay conjectured that the theory flows to the diagonal minimal model $A_{8}$ with $c=14/15$. This cannot be correct since the invertible symmetry of the diagonal minimal model is just $\mathbb{Z}_{2}$ Ruelle:1998zu . In fact, no bosonic minimal model (with any modular invariant) can support a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry, not even as a subgroup (an enhancement to $S_{3}$ is not possible due to Lagrange’s theorem). Fermionized minimal models can have this symmetry, but we must flow to a bosonic theory. Therefore, it must be the case that the relevant CFT to describe this transition has $c\geq 1$.
@@ 2.2 Continuum limit and candidate fixed points
A natural guess would be that the theory flows to the coset Goddard:1984vk ; Goddard:1986ee ; DiFrancesco:1997nk  [^9] We thank Yifan Wang for first suggesting to us this CFT as a candidate.
$\frac{SU(2)_{3}\times SU(2)_{3}}{SU(2)_{6}}\,,$ (34)
@@
which has an invertible center $\mathbb{Z}_{2}$ as well as a complex conjugation $\mathbb{Z}_{2}$ which swaps the two factors in the numerator, along with two Fibonacci line $1_{L,R}$ that satisfy the truncated $SU(2)$ tensor product $1_{L,R}\times 1_{L,R}=0_{L,R}+1_{L,R}$, ensuring the existence of some twofold degeneracies as discussed above. However, this theory has $c=1.35$ and is disfavoured by the numerics as a stable fixed point. Furthermore, it has a relevant operator which is invariant under the full non-invertible symmetry, as we show in Appendix A, and therefore can not be the RG stable fixed point. It is instead a natural candidate for CFT in Figure 8.
A bigger group that is compatible with $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ is $U(1)\times U(1)$ in the $c=1$ compact boson or $D_{8}$ in the $c=1$ orbifold branch. However, the non-invertible symmetries of these conformal manifolds do not seem to exhibit Fibonacci lines at least in any obvious way (see Thorngren:2021yso for a detailed discussion of non-invertible symmetries in $c=1$ CFTs).
For a more systematic analysis, we can list all the known rational values of CFT central charges in the range $1<c\leq 1.2$ (comfortably within our error bars for the IR central charge), for which there can be many possible bosonic theories, related by orbifolds, for example. In this interval the list contains only three values of $c$:

### v3

@@ 1 $\epsilon$ $\sigma$ $\sigma^{\prime}$ $\epsilon^{\prime}$ $\epsilon^{\prime\prime}$
$H=J\sum_{i}H^{(1)}_{i}+J\sum_{i}H^{(2)}_{i}+K\sum_{i}H^{(1)}_{i}H^{(2)}_{i}\,$, (23)
which manifestly preserves two copies of the Fibonacci symmetries. The $K$-term breaks independent one-site translations of the chains, but preserves the diagonal translational symmetry. For $J>0$ and $J<0$, this corresponds to breaking the $\mathbb{Z}_{2}^{(1)}\times\mathbb{Z}_{2}^{(2)}$ and the $\mathbb{S}_{3}^{(1)}\times\mathbb{S}_{3}^{(2)}$ symmetries to their diagonal subgroups respectively. The phase diagram is summarized in Fig. 6. For $J>0$ and $K=0$ we find that the two decoupled tricritical Ising CFTs ($c=1.4$) separate a weakly first order phase transition ($K<0$) in red from a gapped phase ($K>0$). Instead, for $J<0$ and $K=0$ we find a two decoupled 3-state Potts CFT ($c=1.6$), which in fact belongs to a conformal phase for a large range of values of $K$. Making $K$ sufficiently negative ($K/J\approx 1.57$), we eventually find a non-trivial CFT with $c\approx 1.77$ which we refer to as CFT1 and which separates the $c=1.6$ conformal phase from a gapped regime. By making $K$ even more negative ($J/K=0.23$) we find a second non-trivial critical point, denoted CFT2, which has $c=1.35$ and that separates the gapped phase from the weakly first order phase transition. In the remainder of this section we will discuss in detail each of these regions of the phase diagram.
Figure 6: The phase diagram of the 2-coupled anyon chain model. The red region is a pseudo-critical region, which is a weakly first-order phase transition. The blue region is a critical phase described by two decoupled Potts models. The unmarked regions are gapped. The two purple dots are CFTs. We do not know how to identify CFT1 with any rational CFT. CFT2 can be identified as the coset CFT $\frac{SU(2)_{3}\times SU(2)_{3}}{SU(2)_{6}}$
2.1 CFT1
It was noticed in Feiguin et al. (2007) that the single anyon chain model with negative coupling corresponds to the 3-state Potts model. Similarly to the tricritical Ising model, the $\mathbb{Z}_{3}$ symmetry of the 3-state Potts model is realized as a lattice translation by one or two sites. In this case, operators charged under the $\mathbb{Z}_{3}$ symmetry have momentum $\pm\frac{2\pi}{3}$. Therefore, at the J=1J=-1 and $K=0$ point, the 2-coupled anyon chain model (23) corresponds to two decoupled copies of the 3-state Potts CFTs. Turning on the $K$ coupling does not lead to a gapped phase, and the model continues to be described by the decoupled Potts CFT with $c=1.6$. This conformal phase is robust because the CFT that describes it has no relevant operators invariant under
$((\rm{Fibonacci})^{2}\rtimes\mathbb{Z}_{2})\times\mathbb{S}_{3}\,$, (24)
which is the symmetry of the lattice model. The leading irrelevant corrections are described by the action (7).
The critical point denoted as CFT1 in Fig. 6 corresponds to a second order phase transition point separating the critical phase (two decoupled Potts CFTs) and a gapped phase. We can study the effective central charge around this critical point. This is summarized in Fig. 7.
Figure 7: The effective central charge as we vary the couplings near CFT1.
To get the central charge of CFT1, we study the entanglement entropy between a line segment with length $l$ and the result of the spin chain with length LlL-l at the critical point $K/J\approx 1.57$. This is reported in Fig. 8.
Figure 8: The scaling of the entanglement entropy with the size of the cut interval for CFT1. The data points are calculated using lattice size $L=60$ and maximum bond dimension $\chi=5000$.
We fit the effective central charge using the expression
$c_{\rm eff}(l)=c+e^{-2\omega t},\quad{\rm with}\quad t=\log(l_{\rm conformal}),$ (25)
and find
$c_{1}\approx 1.77\,.$ (26)
We justify and discuss the formula (25) in Appendix B. The quality of the fitting can be analyzed using the contour plot of the fitting residues, as shown in Fig. 9.
Figure 9: Left: Fit of the effective central charge using (25). The data points are obtained by doing a linear fit of four neighboring points in Fig. 8 and extracting the slope. The five points near the boundary are discarded because they are likely to be affected by the UV details of the lattice. Right: The contour plot of the fitting error.
To show that this critical point is an actual CFT, we need to check the dynamical critical exponent, which can be calculated by studying how the ground state energy scales with the lattice size. This is summarized in Fig. 10. One can see that the ground state energy follows (9), as expected for a Lorentzian symmetric (zz=1) conformal field theory  [^5] In a general critical theory with $z\neq 1$ one instead expects to observe a scaling $L^{-z}$ for the leading non-extensive part of the ground state energy, see for example Wang and Sedrakyan (2022)..
Figure 10: The ground state energy as a function of the lattice size $L$. This is measured at the critical point CFT1. The linear plot indicates that the system is a CFT with dynamical critical exponent $z=1$.
We can also study the spectrum of this CFT. We have already measured the central charge $c$ in Fig. 8. The result in Fig. 10 then allows us to measure $E_{\infty}$ and the effective “speed of light” $v$ that appears in equation (9). Together they give us the relation between the energy spectrum and the scaling dimension of operators. They are summarized in Fig. 11.
Figure 11: The operator spectrum of CFT1. The red dots/blue circles indicate operators that are even/odd under the permutation symmetry $\mathbb{Z}_{2}^{(p)}$.
Remarkably, we sometimes find degeneracy between states which are even and odd under the $\mathbb{Z}_{2}^{(p)}$ symmetry. The degeneracy is obvious in the decoupled IR theory. Given any two operators $\mathcal{O}$ and $\mathcal{O}^{\prime}$ in a given copy of the Potts model, we can always use them to build two degenerate states in the tensor product theory:
$(\textrm{Fibonacci}\times\textrm{Fibonacci})\rtimes\mathbb{Z}_{2}^{(p)}\,$, (28)
in full analogy with the existence of two dimensional representations for the invertible non-abelian symmetry $D_{4}=(\mathbb{Z}_{2}\times\mathbb{Z}_{2})\rtimes\mathbb{Z}_{2}$, where the rightmost $\mathbb{Z}_{2}$ acts by permuting the actions of the first two. [^6] Consider the two-dimensional Hilbert space spanned $|\sigma\rangle\otimes|1\rangle$ and $|1\rangle\otimes|\sigma\rangle$, the symmetry acts on it as $W^{(1)}=\left(\begin{array}[]{cc}-\varphi^{-1}&0\\ 0&\varphi\\ \end{array}\right),\quad W^{(2)}=\left(\begin{array}[]{cc}\varphi&0\\ 0&-\varphi^{-1}\\ \end{array}\right),\quad\mathbb{Z}_{2}^{(p)}=\left(\begin{array}[]{cc}0&1\\ 1&0\\ \end{array}\right).$ (35) One can easily see that the two-dimensional Hilbert space does not have an invariant subspace under the action of the above matrices. The two states, therefore, should have the same energy. The precise mathematical statement is related to the G-extension of modular tensor categories, see for example Delaney (2019); Benjamin et al. (2025). We will leave a rigorous study of this symmetry to future work. The upshot of this analysis is that CFT1 is a candidate for a UV completion of two decoupled 3-state Potts models whose leading irrelevant interactions are described by (7).
2.2 The pseudo critical phase
We now move our attention to the region near $J=1,K=0$. At the $J=1$, $K=0$ point, the theory becomes two-decoupled anyon chains; the continuum limit can therefore be described as two decoupled tri-critical Ising models. The interaction
$\sum_{i}H^{(1)}_{i}H^{(2)}_{i}\sim\int dx\penalty 10000\ h^{(1)}(x)h^{(2)}(x)\sim\int dp\penalty 10000\ h^{(1)}(-p)h^{(2)}(p)\,$, (36)
@@ 2.2 The pseudo critical phase
The coefficients in front of the continuum operators depend on the details of the lattice model and are unfixed. The $\sigma^{\prime(1)}\sigma^{\prime(2)}$ operator is a relevant operator under the lattice symmetry, which in this case becomes
$((\rm{Fibonacci})^{2}\rtimes\mathbb{Z}_{2})\times\mathbb{Z}_{2}\,.$ (38)
We can study the effective central charge $c_{\textrm{eff}}$ as we vary $K$: the results are summarized in Fig. 12 and suggest that at large negative values of $K/J$, $c_{\textrm{eff}}$ remains constant.
Figure 12: The effective central charge as a function of K/J-K/J. This is a measurement at $L=80$.
We can therefore focus on the $K/J=$-$\infty$ point to discuss the pseudo-critical behavior. The entanglement entropy is reported in Fig. 13.
@@
Picking the data points with $\log(l_{\rm conformal})>3.25$ and fitting against equation (10), we get the central charge
$c_{\rm eff}=1.151\pm 0.001.$ (39)
The error reported here is the standard error of the linear-fit parameter. We can also study the dynamical critical exponent.
@@
As one can see in Fig. 14, the ground state energies follow the scaling relation (9). We can even study the spectrum of this “CFT” like we did for CFT1. The energy eigenvalues are summarized in Fig. 15.
@@
Figure 15: The spectrum of the two coupled chains measured at $K/J=\infty$. The size of the system is $L=40$. The system has a $\mathbb{Z}_{2}^{(p)}$ permutation symmetry which exchanges the states on the two chains. The red dots/blue circles indicate states that are even/odd under the $\mathbb{Z}_{2}^{(p)}$ symmetry. In addition to $\mathbb{Z}_{2}^{(p)}$, the system has an emanant $\mathbb{Z}_{2}$ symmetry, which corresponds to translating the lattice by one site. This can be seen from the fact that the low-energy states cluster around $p=0$ and $p=\pi$.
All these results seem to suggest that we have a critical phase and furthermore that it is a relativistic CFT with dynamical critical exponent $z=1$. What gives away that we are actually in the presence of pseudo-critical behavior is a detailed study of how the effective central charge changes as we measure it for different system sizes, as shown in Fig. 16.
Figure 16: Left: The effective central charge as a function of $l_{\rm conformal}$ for the pseudo critical phase. Right: The contour plot of the residue of the fittings using (25).
Notice that the effective central charge never reaches a constant value, which means we cannot be really be at a critical point. In fact, it is possible to show that the $K/J\to\infty$ limit of our Hamiltonian (23) is integrable and gapped, but with a large correlation length. The authors of Blakeney et al. (2025) were able to map our Hamiltonian to a single integrable anyon chain for the fusion category $\rm{Fib}^{2}$ which exhibits a large correlation length $\xi\approx 155$. The integrability of the Hamiltonian is proven via its connection to the Temperley-Lieb algebra. The Temperley-Lieb algebra also played a key role in previous work in the Euclidean transfer-matrix formalism Fendley and Jacobsen (2008); Vernier et al. (2014) which had also observed this phenomenon: the infinite coupling limit of two $Q$-state Potts models in the loop formulation is described by a single $Q^{2}$-state Potts model, which is gapped for $Q>2$, as is the case for the tricritical Ising model, which has $Q=\varphi\approx 2.62$. [^7] We thank Hubert Saleur for first pointing this out and Jesper Jacobsen for illuminating discussions on this point. In fact, this means that our weakly first order phase transition corresponds to the walking RG behavior in the vicinity of a complex CFT Gorbenko et al. (2018a, b).
Regardless of the weakly first order nature of the phase transition, there is a puzzle in interpreting our numerical result. If we deform the decoupled tricritical Ising model by the term
$g\int d^{2}x\penalty 10000\ \sigma^{\prime(1)}\sigma^{\prime(2)},$ (40)
the physics should not depend on the sign of the coupling, since we can always shift the sign of the coupling by the field redefinition
$\sigma^{\prime(1)}\rightarrow-\sigma^{\prime(1)},$ (41)
which is a symmetry of the undeformed theory. Our lattice model clearly depends on the sign of the coupling constant $K$. This indicates that the irrelevant operators in (37) are dangerously irrelevant operators.
Figure 17: Two possible scenarios of the RG flows of the 2-chain model near $J=1$ and $K=0$. The blue arrow corresponds to the $K/J$. Notice the plots are symmetrical along the vertical axis.
Combining our numerical result with various earlier works, we now have a fairly complete understanding of $\mathcal{C}$-preserving RG flows from the decoupled tricritical Ising fixed point ($2\times\mathcal{M}_{5,4}$). The flows are summarized in Fig. 17. The ($2\times\mathcal{M}_{5,4}$) fixed point has one relevant direction. Depending on the sign of $K/J$, the RG flow can reach either the pseudo-critical phase or the gapped phase. This is conveniently explained by the existence of another CFT, which we now denote as CFT2. In the next section, we will show that CFT2 can also be found in our phase diagram and confirm that there are indeed RG flows from CFT2 to the gapped phase and pseudo-critical phase. However, the precise location of CFT2 in the RG diagram remains obscure. In fact, the pure $\sigma^{\prime(1)}\sigma^{\prime(2)}$ deformation was claimed to lead to a gapped phase in LeClair et al. (1998), even though an explicit construction of the S-matrix was skipped in their work. This supports the second scenario in Fig. 17. This scenario, however, indicates that the lattice model with small K/J-K/J should be in the gapped phase, instead of the pseudo-critical phase. We do not see such a behavior in our DMRG simulation, see Fig. 12. This leaves the possibility that the actual RG flows follow the first scenario in Fig. 17. Notice also that the $\epsilon^{\prime\prime}\equiv\phi_{(3,1)}$ deformation is the leading irrelevant operator that controls the (inverted) Zamolodchikov flows, i.e. we expect to be able to reach a fully unstable decoupled tetracritical Ising fixed point ($2\times\mathcal{M}_{6,5}$) by tuning an additional parameter in the lattice model. [^8] The $\sigma^{\prime(1)}\sigma^{\prime(2)}$ deformation of ($2\times\mathcal{M}_{6,5}$) also leads to a massive phase, this is again discussed in LeClair et al. (1998).
2.3 CFT2: the coset CFT $(SU(2)_{3}\times SU(2)_{3})/SU(2)_{6}$
As we mentioned previously, we need an extra CFT with the symmetry
$((\rm{Fibonacci})^{2}\rtimes\mathbb{Z}_{2})\times\mathbb{Z}_{2}\,$, (42)
@@ 2.3 CFT2: the coset CFT $(SU(2)_{3}\times SU(2)_{3})/SU(2)_{6}$
to explain the RG flow. A natural candidate is the coset Goddard et al. (1985, 1986); Di Francesco et al. (1997) [^9] We thank Yifan Wang for first suggesting to us this CFT as a candidate.
$\frac{SU(2)_{3}\times SU(2)_{3}}{SU(2)_{6}}\,.$ (43)
@@
We provide a brief review of $SU(2)$ coset CFTs in Appendix A. For our analysis, we simply need to note that the coset (43) has an invertible center $\mathbb{Z}_{2}$ symmetry as well as a complex conjugation $\mathbb{Z}_{2}$ which swaps the two factors in the numerator, along with two Fibonacci lines $1_{L,R}$ that satisfy the truncated $SU(2)_{3}$ tensor product $1_{L,R}\times 1_{L,R}=0_{L,R}+1_{L,R}$, ensuring the existence of some twofold degeneracies as discussed above. This theory has $c=1.35$ and furthermore contains a relevant operator with scaling dimension $\Delta=1.5$ which is invariant under the full non-invertible symmetry, as we show in Appendix A.1.
As anticipated above, this CFT also appears in our phase diagram. In Fig. 12, we studied our lattice model (23) up to $K/J=$-$\infty$. This corresponds to fixing $J=0$ and K=1K=-1. We can continue to study the region with $J/K>0$, as shown in Fig. 18.
Figure 18: The effective central charge as a function of ${\rm Arctan}(K/J)$. This is a measurement at $L=80$, with maximum bond dimension $\chi=2000$. The region with ${\rm Arctan}(K/J)<\pi/2$ is the same as what was plotted in Fig. 12. The full vertical line corresponds to $K=0$ while the dashed vertical line corresponds to $K=$-$\infty$. Finally the dashed horizontal line sits at $c_{\textrm{eff}}=1.35$ and is neatly saturated by the point we identify as CFT2.
There exists a local maximum of the effective central charge $c\approx 1.35$ at $J/K\approx 0.23$ which we identify as CFT2. Notice that CFT2 is a critical point separating a pseudo-critical phase and a gapped phase, as predicted in Fig. 17.101010This effective central charge plot closely resembles Figure 5 of Fendley and Jacobsen (2008), which studies a similar model in the euclidean formalism. We can use the standard “data collapsing” method to study the critical exponents of this critical point, as shown in Fig. 19. These turn out to be consistent with the existence of the relevant singlet of dimension $\Delta=1.5$ identified in Appendix A.1.
Figure 19: The “data collapsing” analysis of the critical point corresponding to CFT. Here $E_{0}$ is the ground state energy and $E_{1}$ is the energy of the leading state in the $Z_{2}^{(p)}$ odd sector. We take $1/\nu=0.5$ and $J_{c}/K=0.236$. The blue/orange dots correspond to data measured at $L=40$ and $L=60$, respectively.
3 3-coupled anyon chains
In this section, we will identify an IR fixed point of the three coupled tricritical Ising models using conformal perturbation theory at small but finite coupling. This fixed point was conjectured to be irrational in Chang et al. (2019). Then, using DMRG we identify a stable conformal phase that we conjecture to be described precisely by this putative irrational CFT.
3.1 Naive conformal perturbation theory
We begin by arguing for the existence of a reliable fixed point with $N\geq 3$ coupled tricritical Ising models. We recall the formal action
$S=\sum_{a=1}^{N}S_{\rm tri-Ising}^{(a)}+K\int d^{2}x\sum_{a\neq b}\sigma^{\prime(a)}\sigma^{\prime(b)}\,.$ (44)
To study this model in conformal perturbation theory (CPT), we need to understand the structure of the self-OPE of the perturbation. For the $N=2$ case we have
$\sigma^{\prime(1)}\sigma^{\prime(2)}\times\sigma^{\prime(1)}\sigma^{\prime(2)}\sim 1+(\epsilon^{\prime\prime\,(1)}+\epsilon^{\prime\prime\,(2)})\,$, (45)
which does not contain any relevant operators (and in particular does not contain $\sigma^{\prime(1)}\sigma^{\prime(2)}$ itself). Using the standard result for the beta function in conformal perturbation theory Zamolodchikov (1987)
$\beta_{i}=(2-\Delta_{i})g^{i}-\pi\sum_{j,k}C_{jk}^{i}g^{j}g^{k}+O(g^{3})\,$, (46)
it becomes clear that there is no one-loop fixed point. Note that for a systematic CPT analysis, we need a small parameter $\epsilon$ such that $\Delta=2-\epsilon$ and that UV correlation functions can be computed in an expansion in $\epsilon$. For a small but finite value of epsilon (such as $1/4$ in our case), we can only use the leading order formulas (see for example Komargodski and Simmons-Duffin (2017); Rong and Su (2023) for other examples of such “uncontrolled” conformal perturbation theory). Fortunately, for $N\geq 3$ a one-loop fixed point is actually available. Indeed, noting that
$\sigma^{\prime(i)}\sigma^{\prime(j)}\times\sigma^{\prime(j)}\sigma^{\prime(k)}\sim\sigma^{\prime(i)}\sigma^{\prime(k)}+\textrm{irrelevant}\,$, (47)
allows us to obtain
$\beta_{K}=\frac{K}{4}-\pi\frac{6}{\sqrt{3}^{3}}K^{2}=0\,$, (48)
where 6 is a combinatorial factor for the OPE coefficient and the factors of $\sqrt{3}$ ensure unit normalization of the perturbing operator. This beta function has a UV fixed point at $K=0$ and an IR fixed point at $K^{*}=\sqrt{3}/(8\pi)\approx 0.07$, which is numerically small, suggesting the fixed point might be trustworthy. Once the one-loop fixed point is established, it is straightforward to use Cardy’s c-theorem sum-rule Zamolodchikov (1986); Cardy (1988)
$\Delta c=\frac{\epsilon^{3}}{(C_{KK}^{K})^{2}}\,$ (49)
to obtain an estimate for the IR central charge
$c_{IR}\approx 2.088\,$, (50)
which is remarkably close to the UV value, as is expected from the cubic nature of the central charge correction. Below, we will find a stable conformal phase with a central charge compatible with this value.
3.2 Numerical analysis
We now study the $N=3$ model numerically. We first study the effective central charge as a function of $K/J$. The result for the $K/J>0$ and $K/J<0$ regions are shown in Fig. 20 and Fig. 21 respectively.
Figure 20: The effective central charge of the three-chain model in the $K/J>0$ region. We observe a conformal phase, which corresponds to the proposed irrational conformal field theory. The results are obtained using $L=20$ and maximum bond dimension $\chi=1500$.
Figure 21: The effective central charge of the three chain model in the $K/J<0$ region. This is a gapped phase. The results are obtained using $L=20$ and maximum bond dimension $\chi=1500$.
Since the $N=3$ model has a large local Hilbert space and the corresponding CFTs have large central charges, we can only study the model at fairly small system size $L\lesssim 24$. As before, we calculate the effective central charge by peforming a linear fit of the entanglement entropy as a function of $\log(l_{\rm conformal})$. In practice, we need to drop the points where $l$ is too small and the final results depend on how many data points were dropped, as indicated by the different colors in the figures.
In the $K/J>0$ region, we observe a stable conformal phase. We claim that this corresponds to the irrational CFT discussed in the previous section. Remarkably, as predicted by conformal perturbation theory, the central charge of the IR irrational CFT is very close to the UV CFT. There exists a 2nd order transition from the conformal phase to a gapped phase at $K/J\approx 0.6$. A natural conjecture for this phase transition is that it is described by the CFT consisting of three copies of the tricritical Ising model ($c=2.1$). In the $K/J<0$ region, the effective central decreases sharply, indicating that this is a gapped phase.
We can try to further increase the lattice size to $L=24$, and use higher bond dimension ($\chi=2400$) to study the irrational CFT more carefully. We set $K/J=0.25$, and the result is given in Fig. 22.
Figure 22: The entanglement entropy measure at $K/J=0.25$. The results are calculated using $L=24$ and maximum bond dimension $\chi=2400$.
From the entanglement entropy, we get the central charge
$c=2.10\pm 0.03.$ (51)
This is calculated by performing a linear fit after dropping the five points with the smallest value of $l$. Due to the small lattice size, we are not yet able to perform a more sophisticated finite-size analysis; we will leave this for future work.
Similarly to the $N=2$ chains, we can calculate the dynamical critical exponent of the conformal phase by checking how the ground state energy depends on the lattice size $L$. The result is shown in Fig. 23. This shows that the critical phase is indeed a Lorentzian symmetric conformal field theory.
Figure 23: The ground state energy density $E_{0}/L$ versus the (squared) lattice size $L^{2}$ at the irrational conformal phase. The data is calculated at $K/J=0.25$.
Comments on the Landau-Ginzburg formulation
Let us briefly comment on a possible Landau-Ginzburg realization of the same CFT. As is well known, the diagonal Virasoro minimal models $\mathcal{M}_{m,m+1}$ can also be realized as the IR fixed point of a non-compact scalar with a multi-critical potential Zamolodchikov (1987)
which can be studied in the $3-\epsilon$ expansion Osborn and Stergiou (2018, 2021). It is natural to conjecture that there exists a fixed point for this Lagrangian that can be dimensionally continued to our CFT in 2 dimensions.
4 Comparison to known RCFTs
We now attempt to list candidate RCFTs that could describe fixed points in our phase diagram. While no rigorous classification of RCFTs in increasing order of $c>1$ exists, [^11] An orthogonal classification scheme, where one instead increases the number of chiral primaries has produced interesting results Mathur et al. (1988); Mukhi and Rayhaun (2023), but is not sufficient for our purposes. all known RCFTs can be built as cosets of WZW models or obtained from Narain moduli spaces Yin (2017). We will use two major classes of constructions to generate infinite families of RCFTs. These are cosets of the type
$\frac{g_{k}}{U(1)^{r_{g}}}\,,\qquad\frac{g_{k}\oplus g_{l}}{g_{k+l}}\,$, (54)
where $r_{g}$ denotes the rank of gg. [^12] This list is not claimed to be exhaustive in general, but we believe it is exhaustive in the range $c\leq 1.2$. For example, more general cosets of the type $G_{k}/H_{k}$ with $H$ a subgroup of $G$ such as $G=SU(N)$ and $H=SU(M)$ with $M<N$ also exist but have quickly growing central charges. There are also exceptional conformal embeddings (the so-called Maverick cosets Dunbar and Joshi (1993)) which are believed to exist only for finitely values of the rank, but all known cases have central charges coinciding with the two main families discussed above. We begin by listing the RCFTs in the range $1<c\leq 1.2$. In this interval the list contains only three values of $c$: